OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6. A. y-1=6(x+5) B. y+5=6(x-1) C. y+1=6(x-5) D. y-5=6(x+1)

jimthompson5910 (jim_thompson5910):

y - y1 = m(x-x1) y - (-1) = 6(x-5) y + 1 = 6(x-5)